# Enhancing Multi-Agent Collaboration with Attention-Based Actor-Critic Policies

### v2

@@ IV-C Critic Attention Architecture
IV-D Multi-Agent Baseline
We adopt the approach taken in [15] and set the baseline in the policy gradient by adapting the traditional critic to the multiagent setting in the following way:
$b_{i}(\vec{o}|\vec{a}_{\backslash i})$ $=V(\vec{o}|\vec{a}_{\backslash i})$
$=\sum_{a_{i}\in A_{i}}\pi^{(i)}_{\mathbf{u}}\left(a_{i}\mid\vec{o}\right)\,Q^{(i)}_{\mathbf{w}}(\vec{o},\vec{a}).$
Here, baseline models the multiagent setting in the sense that it estimates the expected value for a specific agent, conditioning on the observations and actions of all other agents.
IV-E Conformity Loss
We further model and encourage collaboration by adjusting the actor loss function to promote diversity of actions among agents, under the assumption that effective collaboration arises when agents work on separate and diverse tasks. [^11] The value of diversity may not hold in all environments, but we found it to be a useful addition for our experiments. We call this adjustment conformity loss. It is computed by measuring the cosine similarity between each unique pair of updated agent-specific embeddings derived from attention layer outputs.
$\mathcal{L}$ $(u_{m_{1}^{\pi}},u_{A})$
$=\theta_{S}\max\left\{\frac{2}{n(n-1)}\sum_{i}\sum_{i^{\prime}\neq i}S_{ii^{\prime}}(u_{m_{1}^{\pi}},u_{A}),\theta_{B}\right\}$
where:
$S_{ii^{\prime}}(u_{m_{1}^{\pi}},u_{A})=\frac{E_{i}(\vec{o}|u_{m_{1}^{\pi}},u_{A})^{\top}E_{i^{\prime}}(\vec{o}|u_{m_{1}^{\pi}},u_{A})}{\|E_{i}(\vec{o}|u_{m_{1}^{\pi}},u_{A})\|\,\|E_{i^{\prime}}(\vec{o}|u_{m_{1}^{\pi}},u_{A})\|}$
and $\theta_{S}$ and $\theta_{B}$ are hyperparameters used to scale and bound the loss. This contribution to overall loss attains high values when agents’ attention layer outputs are very similar (indicating low diversity) and low values when they are not. [^12] In our implementation, we lower bound this loss at a minimum of $\theta_{B}$ to ensure that parameter gradients are directed toward an intermediate level of action diversity during learning.
V Experimentation
In this paper, we compare multiple algorithms against our novel architecture (TAAC) in a simulated soccer environment. In this environment, teams of 3 agents work together to score goals against an opposing team. A goal is scored when the ball is kicked into the opponent’s goal. Each team is composed entirely of agents running the same algorithm to ensure a fair comparison.
V-A Environment Structure
@@
The simulation represents players as solid circles on a 2D plane capable of moving toward the ball and colliding with it. When a player collides with the ball while in mode “kicking”, the ball gains additional velocity. Additionally, the ball collides nearly elastically with the walls, preventing it from leaving the play area. Figure 3 illustrates the simulation environment.
Figure 3: Simulation Environment for Soccer
At each time step, every agent derives an individual, local observation from the global state. In our environment, the state consists of:
@@
  • The relative position vector to the agent’s own goal and opponents goal.
@@
Based on its observation, an agent selects an action. In our environment, an action is defined by five Boolean inputs: move forward, move right, move down, move left, and kick. An agent may choose any combination of these inputs, provided they are not contradictory (e.g., moving left and right simultaneously). Overall, the total number of possible actions is 18.
@@
It is important to note that all rewards (but the distance reward) can be negative, acting as penalties for behaviours such as moving away from the ball, moving the ball away from the opponent’s goal, or being scored on.
Games are configured to last $T$ time steps, and an episode is defined as a sequence of time steps that terminates either when a goal is scored or when the game concludes. As a result, a single game may consist of multiple episodes.
V-B Training Procedure
We employ curriculum learning [23] to progressively build up skills in this complex environment. The training is divided into the following stages:
  1. 1.
    Stage 1: Score goals against an inactive opposing team, with random ball and player spawns.
  2. 2.
    Stage 2: Introduce an opposing team with randomly selected actions, encouraging players to navigate obstacles and opponents. Continuing with random ball and player spawns.
  3. 3.
    Stage 3: Engage in league play, where a team competes against a randomly selected past version of itself while maintaining random spawn positions.
  4. 4.
    Stage 4: Continued league play with fixed starting positions for each team.
V-C Evaluation Metrics
Performance is evaluated using win rates, goal differentials, and ELO ratings [24]. Each team is assigned an Elo rating, which gets updated based on victory, lost or tie and the Elo rating of the opponent. When interpreting Elo ratings, it is important to note that the absolute values are less meaningful than the relative differences between them. Additionally, Goal differential is measured as the difference between the teams’ scores. In addition to performance, we assess collaboration using several key indicators:
@@ V-C Evaluation Metrics
  • Connectivity: This metric measures the number of unobstructed graph-like connections between players, calculated as the ratio of observed connections to the maximum possible connections $\frac{N(N-1)}{2}$. To ensure only meaningful connections are counted, connections only within a range of distance are included. An optimal connectivity level indicates balanced spacing and collaborative positioning.
@@
Using these metrics, we can comprehensively assess both the performance and collaborative behaviour of the teams.
VI Results
Our evaluation compares the performance of our TAAC algorithm against two benchmark algorithms: PPO [25] and MAAC [7]. PPO was chosen as a baseline because of its widespread use and proven robustness across diverse environments, while MAAC was chosen for being state-of-the-art in multi-agent reinforcement learning and closely aligning with our implementation. Additionally, we included a random model to provide a grounded point of reference.
We trained each algorithm for 7 days across over 100,000 games. Note that, due to its low performance and slower learning, MAAC was not advanced to stages 3 or 4 and thus did not gain experience playing against an opponent. We then assembled a league of 20 teams (5 copies of each algorithm) and randomly selected teams to compete in 5,000 games, during which we collected the following data:
Figures 4 to 6 display our performance results, illustrating the ELO ratings, goal differential, and win rates for each algorithm.
Figure 4: Elo rating for each algorithm over the evaluation period.
Figure 5: Win-rate comparison across the algorithms over the evaluation period.
Figure 6: Goal differential Matrix across the algorithms over the evaluation period. Note: The goal differential is calculated as $Score_{team_{1}}-Score_{team_{2}}$; Where a positive value indicates that team 1 is outscoring team 2 by that magnitude, while a negative value indicates the opposite.
Additionally, we present our findings on the collaborative data. Figure 7 illustrates the connectivity score for each model, Figure 8 shows the frequency of possession swaps within a team, and Figure 9 highlights the average pairwise distance among agents on the same team.
Figure 7: Connectivity comparison across the algorithms over the evaluation period.
Figure 8: Possession swaps comparison across the algorithms over the evaluation period.
Figure 9: Average Pairwise distance comparison across the algorithms over the evaluation period.
VII Analysis
Following extensive training of all algorithms, our results indicate that TAAC is the top performer, as evidenced by its superior ELO rating and win rate (see Figures 4 and 5), with PPO closely trailing. In contrast, MAAC significantly underperforms compared to expectations for a state-of-the-art MARL method, possibly due to insufficient training time or suboptimal hyperparameter settings. Additionally, Figure 6 illustrates head-to-head comparisons, highlighting that TAAC outperformed all other algorithms.
When examining collaboration-oriented metrics (Figures 79), the data suggest that TAAC exhibits more coordinated teamwork. For instance, while Figure 9 indicates that PPO agents are more spread out, Figure 7 shows that TAAC’s agents are more closely connected, which may be linked to more advantageous positioning. In addition, the possession swap metrics in Figure 8 imply that TAAC and PPO frequently swapped possession of the ball, potential indicating tactical passing. MAAC demonstrates an incredibly high possessions frequency, potentially indicating frequent passing.
Video recordings further substantiate these findings. The footage confirms that PPO frequently employs a dedicated goalie while the remaining players focus on attacking, explaining their pronounced spatial separation. Additionally, PPO exhibits clustering among its attackers and pursues the ball aggressively, which also explains its lower overall connectivity and relatively high possession frequency In contrast, TAAC consistently maintains a stable offensive formation that only transitions to a defensive position when necessary, which is evident through its connectivity score. TAAC’s dynamic strategy, characterized by frequent ball possession exchanges and timely passes, demonstrates superior tactical coordination. Video recording of MAAC indicates a similar approach to PPO, but instead having the entire team pursue the ball aggressively, this explain their low pairwise distance and connectivity. Furthermore, MAAC’s high possession frequency is also explained by this behaviour, as possession likely rapidly changes between players when the players are grouped together. Additionally, MAAC was consistently outperformed by both PPO and TAAC, as seen in the performance results.
Overall, these results suggest that our novel algorithm demonstrates enhanced collaborative behaviour and improved performance in environments where teamwork is critical. However, due to MAAC’s suboptimal performance and limited training, we plan to investigate its shortcomings further to enable a fairer comparison with state-of-the-art methods. Looking ahead, a logical next step is to evaluate TAAC in alternative environments and scenarios to better understand its limitations. Moreover, selecting an environment where collaboration is more directly tied to performance will allow us to more rigorously test our contribution.
VIII Conclusion
In this work, we introduced TAAC, a multi-agent reinforcement learning algorithm that leverages multi-headed attention in both the actor and critic networks to foster robust collaboration. Our experiments demonstrate that TAAC outperforms competing methods in win rates, goal differentials, and Elo ratings while exhibiting superior coordinated behaviour, as confirmed by game statistics and game recording analysis. Future work will involve a deeper investigation into MAAC to ensure it performs to its full potential and also to explore this new algorithm in different environments to fully test its capabilities.

### v3

@@ IV-C Critic Attention Architecture
V Experiments
(a) Soccer
(b) BoxJump  [boxjump]
(c) LBF [LBF1, LBF2]
Figure 3: Evaluation environments.
We compare multiple algorithms against our novel architecture (TAAC) across the following environments: a novel simulated soccer environment, a cooperative tower-building environment (BoxJump) [boxjump], and Level-Based Foraging (LBF) [LBF1, LBF2]. The soccer environment assesses performance in settings requiring high-level strategy and fine-grained action control; the BoxJump environment assesses performance with high numbers of agents; and the LBF environment evaluates performance on a simple, popular benchmark for cooperative MARL. In the soccer and LBF environments, agents observe the full state, thereby attributing performance improvements solely to coordination rather than to recovering missing state information. By contrast, observations in BoxJump are partial, allowing us to measure agents’ capability to share local information at scale, testing the algorithm’s ability to communicate/coordinate effectively and to gather only crucial information from other agents. Overall, these environments present a diverse set of cooperative multi-agent problems to assess the algorithms’ performance. The soccer environment is fully described in Appendix A, the BoxJump environment is fully described in Appendix B, and the LBF environment is fully described in [LBF1]. Figure 3 depicts the environments.
Our evaluation compares the performance of our TAAC algorithm against two benchmark algorithms: PPO [PPO] and MAAC [MAAC]. PPO was chosen because of its DTDE approach and its widespread use and proven robustness across diverse environments, while MAAC was chosen due to its CTDE approach and because it closely aligns with our implementation. Additionally, in the soccer environment, we included a random model to provide a grounded point of reference.
V-A Training Procedures
Below, we show describe our training methods to ensure that each model is trained to the best potential policy in the BoxJump and soccer environments. Due to LBF’s simplicity, we did not implement any unique strategy for training.
A.1. BoxJump Training Procedure
To gradually guide the models toward the best policy, we employ curriculum learning [curriculumlearningsurvey]. Given that the algorithm is not constrained by the number of agents, we can use curriculum learning to gradually increase the number of agents. To do this, at each episode, we randomly sample a number of agents; this becomes the number of agents used for that episode. Each curriculum stage uses a different agent-count range, as follows:
  1. 1.
    Stage 1: Randomly select from [2, 3, 4, 5] agents for 30,000 episodes.
  2. 2.
    Stage 2: Randomly select from [5, 6, 7, 8] agents for 20,000 episodes.
  3. 3.
    Stage 3: Randomly select from [9, 10, 11, 12] agents for 2,000 episodes.
All algorithms are trained for the same number of episodes in each stage; however, when moving to a new stage, we promote only the highest-reward policy. This procedure enables agents to first understand the reward structure and dynamics of the environment and then develop a collaborative policy.
A.2. Soccer Training Procedure
We also use curriculum learning [curriculumlearningsurvey] to progressively build up skills in the soccer environment. Training is divided into the following stages:
  1. 1.
    Stage 1: Score goals against an opposing team consisting of random agents, with random ball and player spawns.
  2. 2.
    Stage 2: Engage in league play, where a team competes against a randomly selected prior version of itself, while maintaining random spawn positions for ball and players.
  3. 3.
    Stage 3: Continue league play with fixed starting positions for each team.
League play ensures that the learning agent plays against a progressively more diverse and competent set of opponents. Our implementation of league play is a simplified version of [vinyals2019]. For each episode we randomly sample opponents for the learning agent from a pool containing frozen past copies of agents trained using each of the possible algorithms, including an agent choosing actions randomly. During training, every 750 episodes, we add freeze a copy of the learning agent and add it to the pool.
VI Results
VI-A BoxJump - Results
To evaluate BoxJump, we tested each algorithm at each agent count for 1000 episodes and gathered metrics. For each agent count, we only selected the checkpoint of the algorithm which performed best for the agent count. Figure 4 showcases the performance for each algorithm in the BoxJump environment. If TAAC is a superior approach for learning collaboration at scale then we would expect to see its relative performance increase as the number of agents increases. Figure 4 demonstrates this; TAAC-trained agents are likely to achieve a higher max height than alternative approaches at larger numbers of agents in expectation. The TAAC/alternative height distributions are significantly different at 8 agents and above. We provide more results in this environment in Appendix C.1.
Figure 4: BoxJump: This figure showcases the maximum height achieved for each agent count and algorithm during inference. TAAC (ours) achieves more height as the number of agents increases, in expectation. A Kolmogorov-Smirnov Test and a Bootstrap test suggest that the distributions for MAAC and TAAC are significantly different at 8 agents and above.
VI-B LBF - Results
In the LBF environment, we analyzed the performance and sample efficiency of TAAC vs alternatives across a range of different task complexity levels. Each task varies by the number of players (p), the number of food items to be collected (f), and the size of the board. They also vary by whether all agents are required to pick up each food, thus enforcing the need for more collaboration (c). If TAAC is a superior approach for learning collaboration then we would expect to see its relative performance increase as the collaboration becomes more necessary for task completion and as the difficulty of successfully doing it increases. Figure 5 measures the performance for each algorithm on two tasks in the LBF environment. The results suggest that: 1) TAAC has comparable sample efficiency and performance to the other algorithms in a simple task; and 2) TAAC has better performance with comparable sample efficiency in a more complex task which requires more collaboration. We provide results across all tasks in this environment in Appendix C.2.
Figure 5: Top: LBF - Task: 2p-4f-5x5-c. Two players gather four food in a 5x5 grid world; each gathering requires all players. Bottom: LBF - Task: 3p-6f-6x6. Three players gather six food in a 6x6 grid world. Reward is normalized by food levels in both experiments.
VI-C Soccer - Results
Soccer is a more complicated environment which requires a high level of collaboration and strategy to succeed, mainly because it is partially adversarial. To gauge TAAC’s performance in this environment, we trained it using the league play described above and then competed it against the alternative algorithms by randomly sampling opponents trained using those alternatives during evaluation episodes. We did this for 100,000 games and gathered a range of metrics to get an understanding of the performance and collaboration for each algorithm.
Figure 6 and Figure 7 showcase the performance for each algorithm in the Soccer environment. We hope to see TAAC win games against teams whose players are managed by alternative algorithms and do so with collaboration. In Figure 6 we observe that TAAC’s overall win rate is substantially higher than any other approach. Beyond this, anecdotal testing of TAAC against human players in this environment suggests that TAAC-trained agents have human-level performance. Figure 7 suggests that TAAC-managed players successfully pass significantly more than alternative approaches, indicating more collaboration between team members. We provide more results in this environment in Appendix C.3.
Figure 6: Soccer: This figure showcases the win/tie/loss rate for each algorithm against all others (not including itself). It demonstrates TAAC’s better performance when it was paired up head-to-head with the other algorithms.
Figure 7: Soccer: Frequency of possession swaps across algorithms over the evaluation period. Higher values may indicate more intra-team passing and collaboration. We see how TAAC has a marginally higher frequency of possession swaps, indicating more team passing. A Kolmogorov-Smirnov Test and a Bootstrap test suggest that all pairs of distributions are statistically different.
VI-D Discussion
Overall, these results suggest that TAAC has better or comparable performance to the other algorithms in cooperative environments and that this advantage becomes more apparent when the task requires more collaboration and coordination.
This is clearly demonstrated in the BoxJump environment, where TAAC is able to achieve a higher height than the other algorithms as seen in 4. It is also able to do this more quickly with more reward (see Figures 9 and 8 in the appendix). This superiority is related to task complexity in the sense that it grows as the number of agents increases and coordination becomes more difficult; TAAC height is significantly greater than the alternatives above 8 agents (Figure 4) while for 7 agents and below there is no significant difference in performance with MAAC. This showcases how TAAC scales better at coordinating a larger number of agents.
It is also demonstrated in the soccer environment where TAAC has a higher win rate (Figure 6). It also has a higher Elo rating than the other algorithms (see Figure 14 in the appendix). We suspect that this is due to its improved ability to coordinate and collaborate in a highly strategic and competitive environment. This ability is on display in Figure 7, which shows that TAAC more frequently passes the ball than the other algorithms. Note that passing requires clear line of sight between players; Figure 16 in the appendix indicates that TAAC-managed players do not maintain the most passing lanes. Given that TAAC is outscoring and passing more frequently it is likely that it is using its passing lanes more effectively as opposed to MAAC, which has a large amount of passing lanes but very few passes. From our perspective, this suggests that TAAC learns to create and leverage collaborative opportunities dynamically in a selective fashion, only when it is optimal to do so.
We also see the TAAC-managed players are overall closer together on the field than the players managed by alternatives, as seen in Figure 17 in the appendix; in an anecdotal sense, from watching recordings of the algorithms, TAAC-players appear to have a very dynamic strategy, where either the whole team is attacking or the whole team is defending, consistent with its low pairwise distance and connectivity. These recordings also showcase how TAAC is passing the ball more frequently and is more offensive than the other algorithms.
Being more spread out could indicate a stronger sense of roles and responsibilities. A goalie, defenders, and attackers are likely to control different parts of the field and therefore be more spread out. This is consistent with our observations of play; the other algorithms were more spread out and had a more role-based approach, often with dynamic defenders and even goalies; while this approach shows some form of collaboration, it did not seem to be the best strategy for coordinating three agents. Being able to pass the ball around the offense and coordinate an airtight defense seemed to be the best strategy and the one that required the most collaboration.
Finally, across LBF tasks, TAAC’s advantage appears and then disappears as the player count varies. In two-player tasks seen in Figure 5 top (and Figure 10 in the appendix), we see that TAAC performs as well as the other algorithms. As we increase the players to three, TAAC performs marginally better than the other algorithms (see Figures 5 bottom and 11 in the appendix). As we increase the number of agents to four and five, TAAC becomes comparable to the other algorithms again (Figures 12 and 13 in the appendix).
We hypothesize that this is because coordinating two agents to gather food is a relatively low complexity task while coordinating three agents to gather food is more difficult. As team size grows, the LBF task becomes over-provisioned: if a pickup needs $k$ agents, having $n\geq k$ makes accidental coordination likely, limiting the benefit of richer teamwide inter-agent communication and thereby reducing the marginal benefit of TAAC’s collaboration and yielding comparable performance.
Overall, TAAC is a strong performer in complex MARL environments which require a high level of coordination and collaboration. As shown in the results, outside this domain, TAAC has comparable performance to the other algorithms. We also showcase that TAAC is able to scale better with the number of agents than other algorithms.
VII Conclusion
In this work, we introduced TAAC, a multi-agent reinforcement learning algorithm that leverages multi-headed attention in both the actor and critic networks to foster robust collaboration. We evaluated TAAC in a diverse range of cooperative environments, to understand the conditionality which allows TAAC to outperform the other algorithms. Our experiments demonstrate that TAAC outperforms other algorithms in MARL environments which require a high level of coordination and collaboration. Future work could focus on:
  1. 1.
    comparing against a wider range of algorithms in more environments;
  2. 2.
    measuring how performance differs with more attention layers;
  3. 3.
    implementing explainability methods to analyze how the attention mechanism is used by the agents and what information is being shared.
VIII Appendix
Appendix A : Soccer Environment Description
@@ VIII Appendix
The soccer simulation is played on a top-down 2D plane; the players and the ball are represented as solid circles on the plane. Players can exert a force on the ball by colliding with it. When a player collides with the ball while in mode “kicking", the ball gains additional force. Additionally, the ball collides nearly elastically with the walls, preventing it from leaving the play area.
The observations of the agent are composed of:
@@
In our environment, an action is defined by five Boolean inputs: move forward, move right, move down, move left, and kick. An agent may choose any combination of these inputs, provided they are not contradictory (e.g., moving left and right simultaneously). Overall, the total number of possible actions is 18.
@@
It is important to note that all rewards (except for the distance reward) can be negative, acting as penalties for behaviors such as moving away from the ball, moving the ball away from the opponent’s goal, or being scored on. Games are configured to last $T$ time steps, and an episode is defined as a sequence of time steps that terminates either when a goal is scored or when the game concludes. As a result, a single game may consist of multiple episodes.
A.1. Evaluation Metrics
Soccer is a sport where teamwork is not needed to succeed, such as by having a single player who is capable of dribbling and scoring past all of the opposing team. Because of this, in this environment we decided to measure both metrics which indicate performance (winning-capability) as well as collaboration. Performance is evaluated using win rates, goal differentials, and Elo ratings [minka2018trueskill]. Each team is assigned an Elo rating, which gets updated based on victory, loss or tie and the Elo rating of the opponent. When interpreting Elo ratings, it is important to note that the absolute values are less meaningful than the relative differences between them. Additionally, goal differential is measured as the difference between the teams’ scores. Collaboration is measured through:
@@ A.1. Evaluation Metrics
  • Connectivity: This metric measures the number of unobstructed graph-like connections between players, calculated as the ratio of observed connections to the maximum possible connections $\frac{N(N-1)}{2}$. To ensure only meaningful connections are counted, only connections within a range of distances are included. An optimal connectivity level indicates balanced spacing and collaborative positioning. Intuitively this can be thought of as how many unobstructed passing lanes the team has.
@@
Using these metrics, we can comprehensively assess both the performance and collaborative behavior of the teams.
Appendix B : BoxJump Environment Description
In this environment, agents are 2D 1-meter-square objects in a physics environment. They have the ability to move left and right and jump. Their goal is to jump on top of each other to build the tallest tower so as to push the red line as high as possible.
At each time step, every agent derives an individual observation from the global state. In our environment, the state consists of:
  • Horizontal/Vertical Position of agent.
  • Horizontal/Vertical velocity of agent.
  • Distance to closest box in cardinal directions (maximum distance when no box is in a given direction).
  • Boolean of whether the box can jump, or if it is mid-air.
  • An integer representing the number of agents in the environment.
Based on its observation, an agent selects an action, causing a force to be exerted on the box. The agent has three potential actions: move right, move left, and jump, each exerting a force on the box in the given direction. The episode ends if an agent leaves the playable area or if the highest attainable height of $n$ meters is reached.
The reward structure for an individual agent is:
[…]
where $\text{height}_{\text{max}}=\max(\text{height}_{1},\text{height}_{2},\cdots,\text{height}_{n})$ We chose to cube the max height to further incentivize the agents to build the tallest tower and because the difficulty of adding another box given some height is not linear. Intuitively, each new additional height requires the coordination of all agents currently on the tower, making it increasingly difficult to add another box.
Appendix C : Extended Results & Analysis
C.1. BoxJump Extended Results
As well as the results in Figure 4, we also showcase the reward and steps taken to reach a goal height for each algorithm in the BoxJump environment.
Figure 8: BoxJump: This figure showcases the reward achieved for each agent count and algorithm during inference. We can see how TAAC (Ours) achieves a better reward as the number of agents increases. Reward is measured as the cube of the maximum height achieved, so to optimize reward agents need to build the tallest tower as fast as possible and stay in that position for as long as possible. A Kolmogorov-Smirnov Test and a Bootstrap test suggest that all pairs of distributions are statistically different.
Figure 9: BoxJump: This figure showcases the number of steps taken to reach a goal height of $n$ meters where $n$ is the number of agents. We can see how TAAC (Ours) takes fewer steps to achieve its perfect goal and additionally is capable of reaching the goal height at all for higher numbers of agents.
C.2. LBF Extended Results
For LBF, we performed a range of different tasks and gathered the reward through the training process. In Table I, we showcase the reward for each algorithm for each task. We also plot the reward for each algorithm for each task in Figures 5, 10, 11, 12, 13.
TABLE I: LBF tasks: mean $\pm$ 95% confidence interval reward over the last 500 episodes of training
Task MAAC TAAC (Ours) PPO
2p-2f-5x5-c $0.996\pm 0.026$ $0.998\pm 0.018$ $0.985\pm 0.050$
2p-4f-5x5-c $0.998\pm 0.022$ $0.995\pm 0.032$ $0.994\pm 0.033$
3p-3f-6x6 $0.798\pm 0.136$ $\bm{0.925\pm 0.124}$ $0.754\pm 0.138$
3p-6f-6x6 $0.899\pm 0.113$ $\bm{0.955\pm 0.103}$ $0.764\pm 0.132$
4p-8f-8x8 $0.888\pm 0.067$ $\bm{0.902\pm 0.078}$ $0.877\pm 0.074$
5p-10f-10x10 $\bm{0.903\pm 0.052}$ $0.871\pm 0.112$ $0.834\pm 0.114$
Figure 10: LBF - Task: 2p-2f-5x5-c. 2 players gather 2 food in a 5 by 5 grid world; each gathering requires all players. Reward is normalized by food levels.
Figure 11: LBF - Task: 3p-3f-6x6 3 players gather 3 food in a 6 by 6 grid world. Reward is normalized by food levels.
Figure 12: LBF - Task: 4p-8f-8x8 4 players gather 8 food in an 8 by 8 grid world. Reward is normalized by food levels.
Figure 13: LBF - Task: 5p-10f-9x9 5 players gather 10 food in a 9 by 9 grid world. Reward is normalized by food levels.
C.3. Soccer Extended Results
The following are additional results for the soccer environment as described in Appendix A. We gathered these metrics through the evaluation process, and they can be seen in figures 14, 15, 16, 17. You can see the win rate and possession frequency in Figures 6 and 7 in the main results section.
Figure 14: Soccer: Elo ratings for each algorithm over the evaluation period. Elo ratings are a measure of relative skill between two algorithms. TAAC (Ours) achieves a higher elo rating showcasing its superior performance.
Figure 15: Soccer: Goal differential for each algorithm. Each goal differential is the average difference in goals scored between the team on the column and the team on the row. A positive value indicates that the team on the column is outscoring the team on the row, while a negative value indicates the opposite. TAAC (Ours) achieves a positive goal differential against all other algorithms, showcasing its superiority across all other algorithms.
Figure 16: Soccer: Connectivity for each algorithm. As described in Appendix A, connectivity is measured as the ratio of observed connections to the maximum possible connections $\frac{N(N-1)}{2}$. It intuitively measures the number of open passing lanes. This graph showcases that MAAC has the most, followed by TAAC (ours) and PPO. Kolmogorov-Smirnov Test and a Bootstrap test suggest that all pairs of distributions are statistically different.
Figure 17: Soccer: Average pairwise distance for each algorithm. As described in Appendix A, the average pairwise distance is measured as the average distance between all pairs of agents on the same team. TAAC (Ours) achieves a lower average pairwise distance than the other algorithms, showcasing that it is less spread out than the other algorithms. Kolmogorov-Smirnov Test and a Bootstrap test suggest that all pairs of distributions are statistically different.